OpenStudy (anonymous):

simplify: (10a^2)/(7b^2)÷(25a^3)/(14b^3)

OpenStudy (anonymous):

1/245ab^5

OpenStudy (anonymous):

4b/5a?

OpenStudy (anonymous):

1/245ab^5

OpenStudy (mertsj):

\[\frac{10a^2}{7b^2}\times\frac{14b^3}{25a^3}=\frac{4b}{5a}\]

OpenStudy (zepp):

Lol, that equation thingy is killing me :x

OpenStudy (anonymous):

ha lol u guys are so nice!:)

OpenStudy (mertsj):

Keep working on it. It gets easier.

OpenStudy (zepp):

\[\frac{10a^2}{7b^2}*\frac{14b^3}{25a^3}=\frac{10}{7}*\frac{14b}{25a}=\frac{140b}{175a}=\frac{4b}{5a}\] Ah, at last :D

OpenStudy (mertsj):

Good job!!